# algorithmic

### v1

@@ 2.2 Linearized AMP
The AMP algorithm assumes knowledge of the parameters of the model and the corresponding priors. While these can be learned in practice via expectation maximization procedures it is also beneficial to derive spectral algorithms that require fewer assumptions. A standard way toward these is linearization of AMP around its trivial fixed point as done e.g. in Krzakala et al. (2013).
In Appendix C, instead of directly expanding AMP (Algorithm 1) for small mean estimates $\hat{w}^{z},\hat{v}^{z}\ll 1$ which would give an undesirable non-Markovian dependence on past iterates through the Onsager reaction term, we expand the rBP equations (A.14)-(A.17) and then, calculating the appropriate Onsager correction, do the step from linearized rBP to the linearized AMP power-iteration
@@
where the notation without modality index $z$ signifies the stacked vector, $\hat{v}^{t}=\left(\hat{v}^{X,t}_{1},...,\hat{v}^{X,t}_{d},\hat{v}^{Y,t}_{1},...,\hat{v}^{Y,t}_{d}\right)^{T}\in\mathbb{R}^{2d}$ and $\hat{w}^{t}\in\mathbb{R}^{n_{X}+n_{Y}}$. Also we have split the iteration alternating between $v$ and $w$ sectors into two self-contained iterations with block-structured linear operators
$\Gamma_{v}$ $=\left(\begin{matrix}\frac{\lambda_{X}^{2}}{n_{X}}\sigma^{2}_{v^{X}}\sigma^{2}_{w^{X}}S^{T}_{X}S_{X}&\frac{\lambda_{Y}^{2}}{n_{Y}}c_{v}\sigma^{2}_{w^{Y}}S^{T}_{Y}S_{Y}\\ \frac{\lambda_{X}^{2}}{n_{X}}c_{v}\sigma^{2}_{w^{X}}S^{T}_{X}S_{X}&\frac{\lambda_{Y}^{2}}{n_{Y}}\sigma^{2}_{v^{Y}}\sigma^{2}_{w^{Y}}S^{T}_{Y}S_{Y}\\ \end{matrix}\right)$ (18)
$\quad-\mathrm{diag}$
$\Gamma_{w}$ $=\left(\begin{matrix}\frac{\lambda_{X}^{2}}{n_{X}}\sigma^{2}_{v^{X}}\sigma^{2}_{w^{X}}S_{X}S^{T}_{X}&\frac{\lambda_{x}\lambda_{Y}}{\sqrt{n_{X}n_{Y}}}c_{v}\sigma^{2}_{w^{X}}S_{X}S_{Y}^{T}\\ \frac{\lambda_{x}\lambda_{Y}}{\sqrt{n_{X}n_{Y}}}c_{v}\sigma^{2}_{w^{Y}}S_{Y}S_{X}^{T}&\frac{\lambda_{Y}^{2}}{n_{Y}}\sigma^{2}_{v^{Y}}\sigma^{2}_{w^{Y}}S_{Y}S^{T}_{Y}\\ \end{matrix}\right)$
$\quad-\mathrm{diag}$ (19)
where the linear Onsager correction $-\mathrm{diag}$ amounts to setting the diagonal to zero. The form is true for general zero-mean priors and noise channels. For completeness, the pseudo-code for the linearized AMP iteration is given in Appendix C.
@@
How can we relate this linearized AMP algorithm to canonical spectral methods such as PLS? PLS works on the correlation matrix $XY^{T}$ while linearized AMP combines an estimate from the modality itself with an estimate from the other modality. As a consequence, it is clear that PLS will have a sub-optimal recovery threshold for low correlations, since it sees the modalities only through the correlation matrix. Linearized AMP, on the other hand, combines individual and shared information, however it does so optimally for weak recovery while the performance of estimating $v^{z}_{0}$ in the presence of small noise will be sub-optimal, because as seen from (18) the very accurate estimate of $v^{z}_{0}$ based on the individual modality will be corrupted by a correlated but different estimate of $v^{\bar{z}}_{0}$ based on the other modality.
The nonlinear AMP iteration solves this dilemma by reweighting the blocks in the linearization $\Gamma_{v}(\hat{v}^{X},\hat{v}^{Y})$ as the norm of the estimates grows, yielding both optimal sensitivity and performance. As a consequence, even for very small noise, AMP will never converge in a single step, but require at least two steps due to the switch from weak recovery to precise estimation of the latent signal directions.
@@ 2.4 State evolution
By introducing a set of order parameters we now derive the low-dimensional effective dynamics of rBP in the high-dimensional limit, known as state evolution (SE). Since for $d\to\infty$ AMP tracks the dynamics of rBP, the SE is an effective dynamics of AMP as well. Here we sketch the conceptual steps, commenting on a subtlety in applying the Nishimori identity, and give the simplified form arising for Bayes-optimal priors and Gaussian noise channel. The full derivation is detailed in Appendix D.
@@
conventionally referred to as overlaps (or magnetizations) and self-overlaps we can use that due to independence of the messages, node-averaged quantities concentrate to their mean, which is also the mean over the noise disorder. For such self-averaging quantities one can therefore replace the node average by a disorder average. Note that in (21)-(22) we already dropped the target index of the order parameters for this reason. In this way one finds that the quadratic source terms $K$ concentrate to their mean, while the linear source terms $J$ become Gaussian variables. Finally, Bayes-optimality of the priors enables the use of the Nishimori identities (Nishimori, 2001), which yield the simplification $Q^{z}_{w/v}=|M^{z}_{w/v}|$.
Here we wish to make a technical comment why the absolute value appears as a consequence of the $\mathbb{Z}_{2}$ symmetry being spontaneously broken by the random initialization. We believe this clarifies how to deal with this symmetry with respect to the existing literature on state evolution for similar systems, e.g. (Lesieur et al., 2017; Kadmon & Ganguli, 2018). For the Nishimori conditions to hold at all times, initialization of the mean estimators $\hat{w}^{z},\hat{v}^{z}$ must be at zero, consistent with the mean of the prior distribution. Yet zero is a fixed point of the iteration due to symmetry. In practice, AMP is thus initialized with a small random direction, randomly breaking the symmetry and choosing the global signs between $\hat{w}^{z}$ and $\hat{v}^{z}$. Now, in words, the Nishimori identity (Nishimori, 1980; 2001) states that in a quantity averaged both over the posterior distribution, e.g. $P(w|X)$, and the disorder distribution, we can replace one of any iid. sampled variables from the posterior by a variable sampled from the prior distribution, that is
$\mathbb{E}_{w^{0}}\mathbb{E}_{w_{1},w_{2}\sim P(w|X_{w^{0}})}\left[f(w_{1},w_{2},...)\right]$
$=$ $\mathbb{E}_{w^{0}}\mathbb{E}_{w_{1},w_{2}\sim P(w|X_{w^{0}}})\left[f(w^{0},w_{2},...)\right].$ (23)
However, depending on which direction the $\mathbb{Z}_{2}$ symmetry is broken, $\hat{w}^{z}$ and $\hat{v}^{z}$ are in fact estimators of ±$\pm w^{z}$ and ±$\pm v^{z}$. Therefore we need to replace the variable from the posterior, e.g. $\hat{w}^{X}$, by ±$\pm w^{X}$ depending on the sign of the overlap $M_{w}^{X}$. This results in the relation $Q^{z}_{w/v}=|M^{z}_{w/v}|$, restores the symmetry of the SE equations with respect to the sign of the overlaps, see Figure S1, and avoids the obviously erroneous situation of negative $Q_{w/v}^{z}$ that can arise otherwise.
@@
Refer to (D.24)-(D.33) for the form of the SE equations without Bayes-optimal priors and for general noise channels. Depending on the prior, all or part of the expectations in Equations 25 and 24 can be computed analytically, see Section D.2 for Gaussian and Rademacher-Bernoulli priors.
Figure 2: Phase transition of Bayes-optimal recovery (state evolution, blue lines) as a function of the noise strength, compared to AMP, PLS, linearized AMP and informed AMP. Performance is measured as the squared cosine similarity between estimated and ground-truth vectors, e.g. $\textrm{CS}_{w}^{2}=\frac{M_{w}^{2}}{Q_{w}Q_{w^{0}}}$, where the square removes the arbitrary sign of the overlap arising from the $\mathbb{Z}_{2}$ symmetry. a Continuous transition for Gaussian priors on $w^{z}$ and $v^{z}$. Since the two are very close, results are shown here for $w^{z}$, and those for $v^{z}$ in Figure S3. The weak recovery threshold is $\theta_{\mathrm{IT}}=\theta_{\mathrm{alg}}\approx 1.07$. b First-order transition for Rademacher-Bernoulli (sparse) prior on $w^{z}$ with sparsity $\rho_{z}=0.02$ and Gaussian prior on $v^{z}$. Lighter colors refer to $v^{z}$ and darker colors to $w^{z}$. The vertical lines are the algorithmic weak recovery threshold $\theta_{\mathrm{alg}}\approx 0.61$ (green dashed), the information-theoretic threshold $\theta_{\mathrm{IT}}\approx 0.71$ (black dashed) where the upper branch starts dominating the posterior based on the free energy (D.51) , and the spinodal point $\theta_{\mathrm{sp}}\approx 0.72$ (orange dotted). Parameters are for both $z\in\{X,Y\}$: $\alpha_{z}=1,\sigma_{v^{z}}=1,\sqrt{c_{v}}=0.75$, then for panel a $\lambda_{z}=1,\sigma_{w^{z}}=1$, and for panel b $\lambda_{z}=4,\rho_{w^{z}}=0.02$. Each algorithm performance marker is based on one run at size $d=15000$.
2.5 Algorithmic and information-theoretic weak recovery thresholds
@@
Linearizing the SE, by plugging (24) into (25) then expanding for $M_{w}^{z}=\epsilon^{z}\ll 1$, we can assess the stability of the uninformative state at zero overlaps by computing the maximum eigenvalue $\eta_{+}$ of the resulting $2\times 2$ matrix. For zero-mean priors, where consequently the prior overlaps are zero, the algorithmic weak recovery threshold $\theta_{\mathrm{alg}}$ is defined as the smallest signal-to-noise ratio (snr) above which AMP recovers the latent variables better than drawing from the prior. This threshold takes place when $\eta_{+}=1$. Defining the normalized correlation coefficient $\hat{c}_{v}=\frac{c_{v}}{\sigma_{v^{X}}\sigma_{v^{Y}}}$ and the effective snr
@@
where $\hat{\Delta}_{z}$ as defined in (D.11) reduces to $\hat{\Delta}_{z}=\sigma_{\xi^{z}}^{2}$ for the Gaussian channel. The form of (28) arises intuitively when noting that rescaling the model (1,2) by setting the variances $\sigma_{v^{z}},\sigma_{w^{z}},\sigma_{\xi^{z}}\to 1$ corresponds to rescaling $\lambda_{z}\to\lambda_{z}\frac{\sigma_{v^{z}}\sigma_{w^{z}}}{\sigma_{\xi^{z}}}$, and that $\lambda_{z}\sim\alpha^{-\frac{1}{4}}_{z}$ is the scaling of the BBP transition for each single-view matrix (Benaych-Georges & Nadakuditi, 2012). We find that with zero-mean priors the algorithmic weak recovery threshold $\theta_{\mathrm{alg}}$ is given by the condition
11 11 =!${=}\frac{1}{2}\left(\tilde{\lambda}_{X}+\tilde{\lambda}_{Y}+\sqrt{\tilde{\lambda}_{X}^{2}-2(1-2\hat{c}^{4}_{v})\tilde{\lambda}_{X}\tilde{\lambda}_{Y}+\tilde{\lambda}_{Y}^{2}}\right)$ (29)
for general priors and noise channels, assuming they are Bayes-optimal. For symmetric $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}=\tilde{\lambda}$ this reduces to
@@
It is generally conjectured that no polynomial time algorithm can perform better than the $\theta_{\mathrm{alg}}$ of AMP, see Zdeborová & Krzakala (2016).
For some ranges of parameters the so-called first-order phase transitions may appear in the problem as shown in Figure 2 for sparse prior on $w^{z}$. In those cases, the algorithmic threshold for weak recovery may not coincide with the information-theoretic threshold for weak recovery $\theta_{\mathrm{IT}}$. We define $\theta_{\mathrm{IT}}$ as: the smallest snr at which the overlap of the posterior maximum departs from the overlap achieved by the prior. This can be assessed by the Bethe free energy associated to the state evolution given in Equation D.49. Being the negative log of the posterior, the free energy has two minima inside the spinodal regime of a first-order transition. If the lower-overlap branch is uninformative with zero overlaps, $\theta_{\mathrm{IT}}$ is given by the smallest snr where the minimum of the upper-branch becomes deeper than that of the uninformative branch.
@@
Figure 3: Phase transition and comparison of CCA (green ’$\lambda$’) to PLS (gray ’x’) and the Bayes-optimal performance limit for $c_{v}=0.75$. Here the product of cosine similarities of the $w^{z}$ and $v^{z}$ estimates is shown for visual convenience, and the standard deviations across 10 realizations. The threshold of CCA is known analytically from theorem 2.5 of Bykhovskaya & Gorin (2023), and shown here as $\theta_{CCA}$ (green dashed), using the map of notation $\alpha^{X}\to\tau_{K}$, $\alpha^{Y}\to\tau_{M}$ and $\frac{c_{v}^{2}}{(1+\sigma^{2}_{\xi^{X}}/\lambda^{2}_{X})(1+\sigma^{2}_{\xi^{Y}}/\lambda^{2}_{Y})}\to r^{2}$. a Plot as in Figure 2a but with $\alpha^{z}=4$ (other parameters $\lambda^{z}=\sigma_{w}^{z}=\sigma_{v}^{z}=1$ and $d=5000$), since CCA requires $\alpha>1$ so that the covariance matrices $XX^{T}$ and $YY^{T}$ are invertible. The threshold of CCA ($\theta_{\mathrm{CCA}}\approx 0.55$) is considerably lower than that of PLS. b Varying $\alpha^{z}$ instead of $\sigma^{z}_{\xi}$, here for $\lambda^{z}=2$ while $\sigma^{z}_{\xi}=\sigma_{w}^{z}=\sigma_{v}^{z}=1$ and $dn_{z}=5000^{2}$. The threshold of CCA is $\theta_{\mathrm{CCA}}\approx 3.78$ compared to $\theta_{\mathrm{IT}}\approx 0.04$.
3 Numerical results and phase diagram

### v2

@@ 2.2 Linearized AMP
The AMP algorithm assumes knowledge of the parameters of the model and the corresponding priors. While these can be learned in practice via expectation maximization procedures it is also beneficial to derive spectral algorithms that require fewer assumptions. A standard way toward these is linearization of AMP around its trivial fixed point as done e.g. in [44].
In C, instead of directly expanding AMP (Algorithm 1) for small mean estimates $\hat{w}^{z},\hat{v}^{z}\ll 1$ which would give an undesirable non-Markovian dependence on past iterates through the Onsager reaction term, we expand the rBP equations (45)-(48) and then, calculating the appropriate Onsager correction, do the step from linearized rBP to the linearized AMP power-iteration
@@
where the notation without modality index $z$ signifies the stacked vector, $\hat{v}^{t}=\left(\hat{v}^{X,t}_{1},...,\hat{v}^{X,t}_{d},\hat{v}^{Y,t}_{1},...,\hat{v}^{Y,t}_{d}\right)^{T}\in\mathbb{R}^{2d}$ and $\hat{w}^{t}\in\mathbb{R}^{n_{X}+n_{Y}}$. Below we will also use subscripts $S_{X/Y}$ to refer to $S^{X/Y}$ for notational convenience. Now, in (16) and (17) we have split the iteration alternating between $v$ and $w$ sectors into two self-contained iterations with block-structured linear operators
$\Gamma_{v}$ $=\left(\begin{matrix}\frac{\lambda_{X}^{2}}{n_{X}}\sigma^{2}_{v^{X}}\sigma^{2}_{w^{X}}S^{T}_{X}S_{X}&\frac{\lambda_{Y}^{2}}{n_{Y}}c_{v}\sigma^{2}_{w^{Y}}S^{T}_{Y}S_{Y}\\ \frac{\lambda_{X}^{2}}{n_{X}}c_{v}\sigma^{2}_{w^{X}}S^{T}_{X}S_{X}&\frac{\lambda_{Y}^{2}}{n_{Y}}\sigma^{2}_{v^{Y}}\sigma^{2}_{w^{Y}}S^{T}_{Y}S_{Y}\\ \end{matrix}\right)-\mathrm{diag}$ (18)
$\Gamma_{w}$ $=\left(\begin{matrix}\frac{\lambda_{X}^{2}}{n_{X}}\sigma^{2}_{v^{X}}\sigma^{2}_{w^{X}}S_{X}S^{T}_{X}&\frac{\lambda_{x}\lambda_{Y}}{\sqrt{n_{X}n_{Y}}}c_{v}\sigma^{2}_{w^{X}}S_{X}S_{Y}^{T}\\ \frac{\lambda_{x}\lambda_{Y}}{\sqrt{n_{X}n_{Y}}}c_{v}\sigma^{2}_{w^{Y}}S_{Y}S_{X}^{T}&\frac{\lambda_{Y}^{2}}{n_{Y}}\sigma^{2}_{v^{Y}}\sigma^{2}_{w^{Y}}S_{Y}S^{T}_{Y}\\ \end{matrix}\right)-\mathrm{diag}$ (19)
where the linear Onsager correction takes the form $\mathrm{diag}$ amounts to exactly subtracting the diagonal of the operators, so that $\mathrm{diag}(\Gamma_{v/w})=0$. The form is true for general zero-mean priors and noise channels. For completeness, the pseudo-code for the linearized AMP iteration is given in C.
@@
How can we relate this linearized AMP algorithm to canonical spectral methods such as PLS? PLS works on the correlation matrix $XY^{T}$ while linearized AMP combines an estimate from the modality itself with an estimate from the other modality. As a consequence, it is clear that PLS will have a sub-optimal recovery threshold for low correlations, since it sees the modalities only through the correlation matrix. Linearized AMP, on the other hand, combines individual and shared information, however it does so with optimal sensitivity for achieving any nonzero (weak) recovery, while the performance of estimating $v^{z}_{0}$ in the presence of small noise will be sub-optimal, because as seen from (18) the very accurate estimate of $v^{z}_{0}$ based on the individual modality will be corrupted by a correlated but different estimate of $v^{\bar{z}}_{0}$ based on the other modality.
The nonlinear AMP iteration solves this dilemma by reweighting the blocks in the linearization $\Gamma_{v}(\hat{v}^{X},\hat{v}^{Y})$ as the norm of the estimates grows, yielding both optimal sensitivity and performance. As a consequence, even for very small noise, AMP will never converge in a single step, but require at least two steps due to the switch from (weak) recovery to precise estimation of the latent signal directions.
@@ 2.4 State evolution
By introducing a set of order parameters we now derive the low-dimensional effective dynamics of rBP in the high-dimensional limit, known as state evolution (SE). Since for $d\to\infty$ AMP tracks the dynamics of rBP, the SE is an effective dynamics of AMP as well. Here we sketch the conceptual steps, commenting on a subtlety in applying the Nishimori identity, and give the simplified form arising for Bayes-optimal priors and Gaussian noise channel. The full derivation is detailed in D.
@@
conventionally referred to as overlaps (or magnetizations) and self-overlaps we can use that due to independence of the messages, node-averaged quantities concentrate to their mean, which is also the mean over the noise disorder. For such self-averaging quantities one can therefore replace the node average by a disorder average. Note that in (21)-(22) we already dropped the target index of the order parameters for this reason. In this way one finds that the quadratic source terms $K$ concentrate to their mean, while the linear source terms $J$ become Gaussian variables. Finally, Bayes-optimality of the priors enables the use of the Nishimori identities [45], which yield the simplification $Q^{z}_{w/v}=|M^{z}_{w/v}|$.
From here to above (24) we wish to make a technical comment why the absolute value appears as a consequence of the $\mathbb{Z}_{2}$ symmetry being spontaneously broken by the random initialization. We believe this clarifies how to deal with this symmetry with respect to the existing literature on state evolution for similar systems, e.g. [17, 31, 43]. For the Nishimori conditions to hold at all times, initialization of the mean estimators $\hat{w}^{z},\hat{v}^{z}$ must be at zero, consistent with the mean of the prior distribution. Yet zero is a fixed point of the iteration due to symmetry. In practice, AMP is thus initialized with a small random direction, randomly breaking the symmetry and choosing the global signs between $\hat{w}^{z}$ and $\hat{v}^{z}$. Now, in words, the Nishimori identity [46, 45] states that in a quantity averaged both over the posterior distribution, e.g. $P(w|X)$, and the disorder distribution, we can replace one of any iid. sampled variables from the posterior by a variable sampled from the prior distribution, that is
[…]
However, depending on which direction the $\mathbb{Z}_{2}$ symmetry is broken, $\hat{w}^{z}$ and $\hat{v}^{z}$ are in fact estimators of $\pm w^{z}$ and $\pm v^{z}$. Therefore we need to replace the variable from the posterior, e.g. $\hat{w}^{X}$, by $\pm w^{X}$ depending on the sign of the overlap $M_{w}^{X}$. This results in the relation $Q^{z}_{w/v}=|M^{z}_{w/v}|$, restores the symmetry of the SE equations with respect to the sign of the overlaps, see Figure S1, and avoids the obviously erroneous situation of negative $Q_{w/v}^{z}$ that can arise otherwise.
@@
Refer to (94)-(103) for the form of the SE equations without Bayes-optimal priors and for general noise channels. The overall form of these low-dimensional SE equations always follows a generic structure, as expectations over the prior and source term distribution of the denoising function (the mean of the marginal posterior estimate at step $t$) [14, 43]. Depending on the prior, all or part of the expectations in Equations 25 and 24 can be computed analytically, see Section D.2 for Gaussian and Rademacher-Bernoulli priors.
3 Results: Recovery thresholds
3.1 Terminology primer: Algorithmic and information theoretic recovery thresholds
When considering phase transitions in the high-dimensional limit of an estimation problem, multiple distinct thresholds of interest can arise. We focus here on the thresholds in algorithmic and information theoretic recovery, to which we have access through the state evolution equations (24), (25), and an expert reader can skip this section and move to Section 3.2.
Ours is an average-case analysis: By performing a quenched disorder average in the derivation of the SE equations from AMP we compute the threshold for the typical cases exponentially dominating the measure of realizations. This is in contrast to traditional worst-case analyses of algorithms, which consider any exponentially unlikely realizations and can be relevant for safety critical questions far from the high-dimensional limit.
Algorithmic recovery threshold.
We define $\theta_{\mathrm{alg}}$ as the boundary in phase space from which the AMP estimate achieves non-zero overlap with the ground-truth signal. We can access this threshold by a linear stability analysis of the SE equations at the initial (zero) overlap. Such recovery is often called “weak recovery” [47] if the phase transition is continuous, since when passing the threshold the overlap rises continuously from zero and is still small at first.
It is generally conjectured that no (robust to noise) polynomial time algorithm can perform better than BP in the high-dimensional limit, see [13]. Since AMP becomes an exact approximation of BP in this limit for our model, this justifies to treat the algorithmic threshold identified by $\theta_{\mathrm{alg}}=\theta_{\mathrm{AMP}}$ as the optimal value achievable for practical algorithms.
Information theoretic recovery threshold.
We define $\theta_{\mathrm{IT}}$ as the boundary in phase space from which a sample taken from the posterior distribution achieves non-zero overlap with the ground-truth signal (in the exponentially dominating typical cases). By definition, it is information theoretically not possible to achieve a higher performance.
In the presence of a continuous phase transition, $\theta_{\mathrm{IT}}$ here coincides with $\theta_{\mathrm{alg}}$ since the uninformative initial state at zero overlap loses stability and AMP iteration reaches the informative branch dominating the posterior. However, for some ranges of parameters first-order phase transitions may appear in the problem as shown in Figure 3 for sparse prior on $w^{z}$. In those cases, the algorithmic threshold may not coincide with the information-theoretic threshold $\theta_{\mathrm{IT}}$, since the uninformative branch at zero overlap remains a stable minimum, while a second branch with non-zero overlap may actually dominate the posterior. This can be assessed by the Bethe free-energy associated to the state evolution given in Equation 119. Being the negative log of the posterior, the free energy has two minima inside the spinodal regime of a first-order transition, one for each branch. If the lower-overlap branch is uninformative with zero overlaps, $\theta_{\mathrm{IT}}$ is given by the boundary in phase space where the minimum of the informative branch becomes deeper than that of the uninformative branch. Then the posterior is dominated by estimates with positive overlap, so recovery even with substantial overlap is theoretically possible - but algorithms starting with close to zero overlap are trapped in the uninformative minimum for exponentially large times. Such a regime is termed an algorithmically “hard” phase [13], see also [48] for a detailed discussion.
In both cases, note that we are here concerned with thresholds for “recovery” and not “detection”, since the possibility to detect that a signal is present in the data does not generally imply the possibility to find an estimate that correlates with the signal (recovering it), see e.g. [49].
Finally, the term “strong recovery” has then been associated with the jump from zero (or low) correlation to high correlation in a first-order transition. But we do not use this less specific term here because it is possible to construct settings where the jump is small or in which multiple first-order transitions appear [50], while $\theta_{\mathrm{alg}}$ and $\theta_{\mathrm{IT}}$ always remain unique.
3.2 Algorithmic recovery thresholds
@@
Linearizing the SE, by plugging (24) into (25) then expanding for $M_{w}^{z}=\epsilon^{z}\ll 1$, we can assess the stability of the uninformative state at zero overlaps by computing the maximum eigenvalue $\eta_{+}$ of the resulting $2\times 2$ matrix. For zero-mean priors, where consequently the prior overlaps are zero, the algorithmic recovery threshold $\theta_{\mathrm{alg}}$ thus takes place when crossing $\eta_{+}=1$. Note that by construction the threshold coincides for AMP and linearized AMP.
For compact notation we define the normalized correlation coefficient $\hat{c}_{v}=\frac{c_{v}}{\sigma_{v^{X}}\sigma_{v^{Y}}}$ and the effective signal-to-noise ratio (SNR)
@@
[…]
$1$ $\overset{!}{=}\frac{1}{2}\left(\tilde{\lambda}_{X}+\tilde{\lambda}_{Y}+\sqrt{\tilde{\lambda}_{X}^{2}-2(1-2\hat{c}^{4}_{v})\tilde{\lambda}_{X}\tilde{\lambda}_{Y}+\tilde{\lambda}_{Y}^{2}}\right)$ (29)
for general priors and noise channels, assuming they are Bayes-optimal (that is, matching the ground-truth prior and noise channel). The phase diagram based on this threshold is shown in Figure 2 and wi11 be discussed in the next section. For symmetric $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}=\tilde{\lambda}$ this reduces to
@@
In Figure 2 we plot phase diagrams illustrating the algorithmic recovery threshold $\theta_{\mathrm{alg}}$ given through (29) in the reduced three dimensional parameter space of effective SNR’s $\tilde{\lambda}_{X},\tilde{\lambda}_{Y}$ and correlation coefficient $\hat{c}_{v}$. Note that $\theta_{\mathrm{IT}}$ may vary depending on the prior and is not shown, while $\theta_{\mathrm{alg}}$ is given by (29) for any zero-mean prior. Figure 2a shows for $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}$ the interpolation between zero correlation, equivalent to two single-view models, and perfect correlation, equivalent to the stackable model (20). Figure 2b illustrates the improvement of the multi-modal threshold over the thresholds of the two isolated single-view models (dashed black lines). Due to the definition of the SNR (28), this plot can for example be interpreted as independently varying the aspect ratios $\alpha_{z}$. Apart from the gain in the lower left sector where no recovery is possible in any isolated model, note that also in the lower-right and upper-left sectors some degree of recovery is always possible in both modalities when the correlation is nonzero, see also Figure S2.
Figure 2: Phase diagram of the algorithmic recovery threshold based on (29) and (30). a Varying the correlation coefficient for symmetric SNR’s $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}=\tilde{\lambda}$. b The $\tilde{\lambda}_{X},\tilde{\lambda}_{Y}$ plane for $\hat{c}_{v}^{4}=1/2$. Dashed black lines show the thresholds of the modalities in isolation. The dotted line in a and b indicates the intersection of both planes.
Figure 3: Phase transition of Bayes-optimal recovery (state evolution, blue lines) as a function of the noise strength, compared to AMP, PLS, linearized AMP and informed AMP. Performance is measured as the squared cosine similarity between estimated and ground-truth vectors, e.g. $\textrm{CS}_{w}^{2}=\frac{M_{w}^{2}}{Q_{w}Q_{w^{0}}}$, where the square removes the arbitrary sign of the overlap arising from the $\mathbb{Z}_{2}$ symmetry. To reduce clutter, results shown here are for $w^{z}$, and those showing the same qualitative behavior for $v^{z}$ can be found in Figure S3. a Continuous transition for Gaussian priors on $w^{z}$ and $v^{z}$. The recovery threshold is $\theta_{\mathrm{IT}}=\theta_{\mathrm{alg}}\approx 1.07$. b First-order transition for Rademacher-Bernoulli (sparse) prior on $w^{z}$ with sparsity $\rho_{z}=0.02$ and Gaussian prior on $v^{z}$. The vertical lines are the algorithmic recovery threshold $\theta_{\mathrm{alg}}\approx 0.61$ (green dashed), the information-theoretic threshold $\theta_{\mathrm{IT}}\approx 0.71$ (black dashed) where the upper branch starts dominating the posterior based on the free energy (121) , and the spinodal point $\theta_{\mathrm{sp}}\approx 0.72$ (orange dotted). Parameters are for both $z\in\{X,Y\}$: $\alpha_{z}=1,\sigma_{v^{z}}=1,\sqrt{c_{v}}=0.75$, then for panel a $\lambda_{z}=1,\sigma_{w^{z}}=1$, and for panel b $\lambda_{z}=4,\rho_{w^{z}}=0.02$. Each algorithm performance marker is based on one run at size $d=15000$.
@@
Figure 4: Phase transition and comparison of CCA (green $\lambda$’) to PLS (gray ’x’) and the Bayes-optimal performance limit for $c_{v}=0.75$. Here the product of cosine similarities of the $w^{z}$ and $v^{z}$ estimates is shown for visual convenience, and the standard deviations across 10 realizations. The threshold of CCA is known analytically from theorem 2.5 of [22], and shown here as $\theta_{CCA}$ (green dashed). The threshold condition derived by [22] is $r^{2}\sqrt{(\tau_{K}-1)(\tau_{M}-1)}=1$ and we use the map of notation $\alpha^{X}\to\tau_{K}$, $\alpha^{Y}\to\tau_{M}$ and $\frac{c_{v}^{2}}{(1+\sigma^{2}_{\xi^{X}}/\lambda^{2}_{X})(1+\sigma^{2}_{\xi^{Y}}/\lambda^{2}_{Y})}\to r^{2}$. a Plot as in Figure 3a but with $\alpha^{z}=4$ (other parameters $\lambda^{z}=\sigma_{w}^{z}=\sigma_{v}^{z}=1$ and $d=5000$), since CCA requires $\alpha>1$ so that the covariance matrices $XX^{T}$ and $YY^{T}$ are invertible. The threshold of CCA ($\theta_{\mathrm{CCA}}\approx 0.55$) is considerably lower than that of PLS. b Varying $\alpha^{z}$ instead of $\sigma^{z}_{\xi}$, here for $\lambda^{z}=2$ while $\sigma^{z}_{\xi}=\sigma_{w}^{z}=\sigma_{v}^{z}=1$ and $dn_{z}=5000^{2}$. The threshold of CCA is $\theta_{\mathrm{CCA}}\approx 3.78$ compared to $\theta_{\mathrm{IT}}\approx 0.04$.
4 Numerical results and comparison to PLS and CCA